A vector of magnitude 10 has its rectangular components as 8 and 6 along x and y axes.Find the angles it makes with these axes

Grade:11

1 Answers

Sagar Gupta
6 years ago
The vector is having the x-component as 8 and y-component as 6.So It`s angle with the axis will be Tan (Theta) = y-component/x- component = 6/8=3/4 so (Theta) = Tan-1 (3/4)